Using the chemical equation: 2 Ca3(PO4)2 + 6SiO2 + 10C -> 6CaSiO3 + P4 + 10CO. Calculate the minimum mass of Ca3(PO4)2 required to produce 69.8 grams of P4.

1
Expert's answer
2021-11-29T02:36:03-0500

Molar mass of P4= 124

69.8/124=0.563

Molar mass of Ca3(PO4)2= 310.18

= 310.18×2×0.563=349.203g